# Talking to DINO: Bridging Self-Supervised Vision Backbones with Language for Open-Vocabulary Segmentation

### v1

@@ 1 Introduction
To close the gap between vision-language and self-supervised embedding spaces, we propose Talk2DINO, a method that combines the spatial sensitivity of DINOv2 with the text-image alignment capabilities of CLIP, enabling a highly localized, multimodal image understanding. Our approach, depicted in Fig. 1, learns a mapping function that translates the text embeddings of CLIP to interact with DINOv2’s patch-level embeddings without fine-tuning the underlying backbones. In particular, our alignment mechanism enhances text-image correspondence by exploiting self-attention heads of DINOv2 to highlight diverse semantic regions of the image. During training, we weight visual patch features using DINOv2’s attention maps, dynamically selecting the head that best aligns with the provided caption. This visual feature is then used to maximize similarity with the caption through contrastive learning. At inference, we calculate the similarity of visual patches to each textual label, including a novel background cleaning procedure that re-weights class scores based on the self-attention maps.
Our approach demonstrates state-of-the-art performance on several unsupervised OVS datasets by only learning a minimal set of parameters. Our results highlight that a self-supervised vision-only encoder can generate embeddings with semantic properties akin to textual representations, opening up new pathways for addressing spatial understanding limitations in CLIP-like models [40].
To sum up, our main contributions are as follows:
  • We propose Talk2DINO, the first model that directly aligns DINOv2 and CLIP feature spaces for OVS. By mapping CLIP’s textual embeddings into the DINOv2 space using a non-linear warping function, Talk2DINO effectively provides language properties to DINOv2.
@@
  • We showcase the capabilities of Talk2DINO on unsupervised OVS by devising a computationally efficient inference pipeline that also employs a novel approach based on DINOv2 self-attention to clean background objects.
@@ 2 Related Work
Vision-Language Pre-Training. In the last years, vision-language pre-training has gained increasing interest by learning multimodal representations that can be easily transferred to downstream tasks [16, 34, 21]. The popular CLIP model [34] is trained on large-scale web-scraped data through a contrastive objective, which maximizes the similarity of the representations of the corresponding image-text pair while minimizing the similarity of the other pairs within a batch. This approach has demonstrated remarkable zero-shot classification and retrieval performance. However, learning a multimodal representation by matching global images and texts poses challenges in localizing regions with their corresponding text, showing limited performance in tasks based on dense predictions [31, 35, 55, 4].
@@
On the contrary, another research direction investigates how to force the segmentation capabilities to emerge without relying on direct supervision from segmentation data. Cha et al. [9] identify this setting as unsupervised OVS and propose a unified evaluation protocol. Approaches in this line can be categorized into two main groups: (i) training-free methods that propose architectural adaptations to enable pre-trained models to produce localized multimodal features [58, 23, 46, 5, 41, 42, 19], and (ii) weakly-supervised methods that leverage large sets of image-text pairs with dedicated learning strategies that aim to improve the correspondence between regions and texts [49, 9, 35, 50, 28, 36]. Our model lies in the former category since we exploit a weak supervision to learn how to bridge the CLIP and DINOv2 feature spaces.
@@
Self-Supervised Backbones. Recent advances in self-supervised learning have led to models that showcase impressive matching and localization capabilities. In particular, the DINO family of models [8, 33, 12] employs Vision Transformers [14] trained with a self-distillation loss function, and has shown that patch-level features learned through self-supervision can yield semantic information. Moreover, a strong relationship has been observed between self-attention activations and foreground regions of the image. Hence, researchers have recently focused on the usage of DINO for the unsupervised object segmentation task [38, 39, 45, 44, 43].
@@
While the aforementioned models are purely visual, recent works have also focused on connecting self-supervised feature spaces with textual representations for OVS. ReCo [37], OVDiff [22], FOSSIL [2], and FreeDA [3] are training-free approaches that build prototypes in the visual space according to pre-defined textual categories. CLIP-DINOiser [47] demonstrates that CLIP can be fine-tuned with the supervision of DINO to retain improved localization capabilities. ProxyCLIP [23] proposes a proxy attention module to integrate DINO features with values from the last attention layer of CLIP, thus employing two visual backbones at prediction time. Our method is closely related to this research field since we aim to combine the multimodal understanding capabilities of CLIP with the semantic localization properties of DINOv2. However, in contrast to these methods, we propose to directly map the textual representations from the textual encoder of CLIP to the DINOv2 space, and demonstrate that our approach sets a new state of the art without relying on multiple visual backbones or on external sources of knowledge.
@@
Figure 2: Overview of the training methodology of Talk2DINO. We learn a projection $\psi(\cdot)$ that maps the CLIP textual embeddings to the visual embedding space of DINOv2. Given the dense feature map and the attention maps extracted from DINOv2, we generate $N$ visual embeddings by computing a weighted average of the feature map with each attention map. We then compute the similarity between each visual embedding and the projected text embedding, and apply a selection head function to produce a global similarity score. We use the maximum similarity score to align textual and visual spaces.
@@ 3.1 Preliminaries
Conversely, purely visual self-supervised backbones like DINOv2 [33] have shown remarkable semantic and local consistency of spatial embeddings, enabling agnostic image segmentation [39, 38, 45]. These abilities occur naturally in the last attention layer of DINOv2, where the attention maps computed between the CLS token $v^{*}\in\mathbb{R}^{D_{v}}$ and the spatial tokens $v$ align with relevant objects within the image (see Fig. 2). Despite the remarkable results observed on image-only tasks, DINOv2 lacks a solid bridge with natural language, making it impossible to directly replace $v^{*}$ with a text feature $t$ to compute the similarities expressed in Eq. 1.
While DINOv2 and CLIP embedding spaces are traditionally thought as being uncorrelated spaces, we show that the CLIP textual embedding space can be projected into the DINOv2 space through a learnable nonlinear warping followed by an affine transformation.
@@ 3.2 Augmenting DINO with Semantics
Specifically, we first extract the $N$ attention maps $A_{i}\in\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}}$ (one for each of the $i=1,...,N$ heads) which DINOv2 computes between the $v^{*}$ CLS vector and its patch features $v$ from the last layer. One of the key features of DINOv2 is that each $A_{i}$ highlights different semantic regions within the image. For each attention map $A_{i}$, we compute a visual embedding $v^{A_{i}}\in\mathbb{R}^{D_{v}}$ as a weighted average of the dense feature map $v$, emphasizing the spatial areas that $A_{i}$ highlights. We then calculate the cosine similarity between each $v^{A_{i}}$ and the projected text embedding $\psi(t)$, resulting in $N$ similarity scores. Formally, the cosine similarity score between a head and the text embedding is defined as
$\text{sim}(v^{A_{i}},t)$ $=\frac{v^{A_{i}}\cdot\psi(t)^{\intercal}}{||v^{A_{i}}||\,||\psi(t)||},$ (4)
with $\text{with }v^{A_{i}}$ $=\sum_{h,w}v_{[h,w]}\text{softmax}(A_{i})_{[h,w]}.$ (5)
To obtain the most relevant score for alignment, we apply a selection function over the similarity scores obtained for different heads. In particular, we choose the maximum similarity $\max_{i=1,...,N}\text{sim}(A_{i},v,t)$ score across all heads, therefore promoting a robust alignment between textual and visual representations that adapts to the most salient visual features corresponding to the text query.
@@
$\mathcal{L}_{\text{InfoNCE}}=-\frac{1}{B}\sum_{i=1}^{B}\log\frac{\exp(\text{sim}(\psi(t_{i}),\tilde{v}_{i}))}{\sum_{j=1}^{B}\exp(\text{sim}(\psi(t_{i}),\tilde{v}_{j}))}.$ (7)
@@ 3.3 Identifying Background Regions
and normalize $\mathcal{F}$ by applying a softmax normalization over both spatial axes, and linearly re-projecting its values in the range $[\min_{j,h,w}\mathcal{S}(I,T_{j})_{[h,w]},\max_{j,h,w}\mathcal{S}(I,T_{j})_{[h,w]}]$, where $\mathcal{S}(\cdot)$ is the similarity map defined in Eq. 1. We exploit the resulting normalized average attention per category to shape the similarity map by activating the foreground region and deactivating the background. The resulting shaped similarity map $\overline{\mathcal{S}}\in\mathbb{R}^{H\times W}$ is defined as
@@
where $\lambda$ is a hyperparameter representing the relevance of the background shaping in computing the segmentation masks. The segmentation mask for the background region is then simply identified as the collection of pixels for which the shaped similarity map is lower than a threshold across all semantic categories.
@@
Model Visual Backbone Frozen VOC Context Stuff Cityscapes ADE Avg VOC Context Stuff Cityscapes ADE Avg
@@
GroupViT [49] Custom ViT 79.7 23.4 15.3 11.1 9.2 27.7 - - - - - -
ReCo [37] CLIP 57.7 22.3 14.8 21.1 11.2 25.4 - - - - - -
TCL [9] CLIP 77.5 30.3 19.6 23.1 14.9 33.1 - - - - - -
SILC [32] Custom ViT 77.5 31.6 20.8 26.9 19.3 35.2 - - - - - -
MaskCLIP [58] CLIP 72.1 25.3 15.1 11.2 9.0 26.5 29.4 12.4 8.8 11.5 7.2 13.9
CLIP-DIY [46] CLIP+DINO 79.7 19.8 13.3 11.6 9.9 26.9 - - - - - -
OVDiff [22] CLIP+DINO+SD 80.9 32.9 20.3 23.4 14.9 34.5 - - - - - -
SCLIP [42] CLIP 80.4 34.2 22.4 32.2 16.1 37.1 70.6 25.2 17.6 21.3 10.9 29.1
CLIP-DINOiser [47] CLIP 80.9 31.7 24.6 35.9 20.0 38.6 - - - - - -
NACLIP [19] CLIP 79.7 35.2 23.3 35.5 17.4 38.2 78.7 32.1 21.4 31.4 17.3 36.2
ProxyCLIP [23] CLIP+DINOv2 83.0 37.2 25.4 33.9 19.7 39.8 85.2 36.2 24.6 35.2 21.6 40.6
ProxyCLIP [23] CLIP+DINO 80.3 39.1 26.5 38.1 20.2 40.8 83.2 37.7 25.6 40.1 22.6 41.8
Talk2DINO (Ours) DINOv2 85.3 40.5 27.9 38.2 21.8 42.7 86.8 39.3 27.1 37.0 21.4 42.3
@@
GroupViT [49] Custom ViT 81.5 23.8 15.4 11.6 9.4 28.3 - - - - - -
ReCo [37] CLIP 62.4 24.7 16.3 22.8 12.4 27.7 - - - - - -
TCL [9] CLIP 83.2 33.9 22.4 24.0 17.1 36.1 - - - - - -
MaskCLIP [58] CLIP 74.9 26.4 16.4 12.6 9.8 28.0 32.7 14.6 10.5 13.2 8.5 15.9
FOSSIL [2] DINOv2 - - - - - - - 35.8 24.8 23.2 18.8 -
SCLIP [42] CLIP 83.5 36.1 23.9 34.1 17.8 39.1 76.3 27.4 18.7 23.9 11.8 31.6
NACLIP [19] CLIP 83.0 38.4 25.7 38.3 19.1 40.9 84.5 36.4 24.6 37.1 19.6 40.4
FreeDA [3] DINOv2 75.6 37.6 25.2 34.4 20.7 38.7 70.2 39.0 26.9 33.2 19.5 37.8
FreeDA [3] CLIP+DINOv2 85.6 43.1 27.8 36.7 22.4 43.1 87.9 43.5 28.8 36.7 23.2 44.0
Talk2DINO (Ours) DINOv2 87.0 43.5 30.3 40.8 23.5 45.0 88.7 43.2 30.0 39.3 23.4 44.9
Table 1: Comparison with unsupervised OVS models on Pascal VOC [15], Pascal Context [30], COCO Stuff [7], Cityscapes [11], and ADE20K [56, 57]. For each method, we specify the visual backbone used, along with whether it is frozen or fine-tuned.

### v3

@@ 1 Introduction
To close the gap between vision-language and self-supervised embedding spaces, we propose Talk2DINO, a method that combines the spatial sensitivity of DINOv2 with the text-image alignment capabilities of CLIP, enabling a highly localized multimodal image understanding. Our approach, depicted in Fig. 1, learns a mapping function that translates the text embeddings of CLIP to interact with the patch-level embeddings of DINOv2 without fine-tuning the underlying backbones. Our alignment mechanism enhances text-image correspondence by exploiting the self-attention heads of DINOv2 to highlight diverse regions of the image. During training, we weight visual patch embeddings using the attention maps of DINOv2, dynamically selecting the head that best aligns with the provided caption. This embedding is then used to maximize similarity with the caption representation through contrastive learning. At inference time, we calculate the similarity of visual patches to each textual label, including a novel background cleaning procedure that weights class scores using attention maps.
Our approach demonstrates state-of-the-art performance in unsupervised OVS with minimal parameter learning. Our results show that a self-supervised vision-only encoder can generate embeddings with semantic properties akin to textual representations, opening up new pathways for addressing spatial understanding limitations in CLIP-like models [45]. To sum up, our main contributions are as follows:
  • We propose Talk2DINO, the first model that provides language properties to DINOv2 by mapping the CLIP textual embeddings into the DINOv2 space through a non-linear warping function.
@@
  • We showcase the capabilities of Talk2DINO on unsupervised OVS by devising a computationally efficient inference pipeline that also employs a novel approach based on DINOv2 self-attention to improve distinguishing foreground categories from the background.
@@ 2 Related Work
Vision-Language Pre-Training. In the last years, vision-language pre-training has gained increasing interest by learning multimodal representations that can be easily transferred to downstream tasks [16, 20]. The popular CLIP model [38] is trained on large-scale web-scraped data through a contrastive objective, which maximizes the similarity of the representations of corresponding image-text pairs while minimizing the similarity of the other pairs within a batch. This approach has demonstrated remarkable zero-shot classification and retrieval performance. However, learning a multimodal representation by matching global images and texts poses challenges in localizing regions with their corresponding text, showing limited performance in dense prediction tasks [35, 39, 60, 4, 29].
@@
On the contrary, another research direction investigates how to force the segmentation capabilities to emerge without relying on direct supervision from segmentation data. Cha et al[9] identify this setting as unsupervised OVS and propose a unified evaluation protocol. Approaches in this line can be categorized into two main groups: (i) training-free methods that propose architectural adaptations to enable pre-trained models to produce localized multimodal features [63, 26, 51, 5, 46, 47, 18, 25, 44], and (ii) weakly-supervised methods that leverage large sets of image-text pairs with dedicated learning strategies that aim to improve the correspondence between regions and texts [54, 9, 39, 55, 32, 40]. Our model lies in the latter category since we exploit a weak supervision to learn how to bridge the CLIP and DINOv2 feature spaces.
@@
Self-Supervised Backbones. Recent advances in self-supervised learning have led to models that showcase impressive matching and localization capabilities. In particular, the DINO family of models [8, 37, 12] employs Vision Transformers [14] trained with self-distillation, and has shown that patch-level features learned through self-supervision can yield semantic information. Also, a strong relationship has been observed between self-attention activations and foreground regions of the image. Hence, researchers have recently focused on the usage of DINO for the unsupervised segmentation task [42, 43, 50, 49, 48].
@@
While the aforementioned models are purely visual, recent works have also focused on connecting self-supervised feature spaces with textual representations for OVS. ReCo [41], OVDiff [23], FOSSIL [2], and FreeDA [3] are training-free approaches that build prototypes in the visual space according to pre-defined textual categories. CLIP-DINOiser [52] demonstrates that CLIP can be fine-tuned with the supervision of DINO to retain improved localization capabilities. LaVG [22] employs DINO to propose class-agnostic regions and computes the average embedding from CLIP for each region. ProxyCLIP [26] proposes a proxy attention module to integrate DINO features with values from the last attention layer of CLIP, thus employing the two visual backbones at prediction time. Our method is closely related to this research field since we aim to combine the multimodal understanding capabilities of CLIP with the localization properties of DINOv2. However, in contrast to these methods, we propose to directly map the textual representations from the textual encoder of CLIP to the DINOv2 space, and demonstrate that our approach sets a new state-of-the-art without relying on multiple visual backbones or on external sources of knowledge.
@@
Figure 2: Overview of the training methodology of Talk2DINO. We learn a projection $\psi(\cdot)$ that maps the CLIP textual embeddings to the visual embedding space of DINOv2. Given the dense feature map and the attention maps extracted from DINOv2, we generate NN visual embeddings by computing a weighted average of the feature map with each attention map. We then compute the similarity between each visual embedding and the projected text embedding, and use the maximum similarity as the global alignment score.
@@ 3.1 Preliminaries
Conversely, purely visual self-supervised backbones like DINOv2 [37] have shown remarkable semantic and local consistency of spatial embeddings, enabling agnostic image segmentation [43, 42, 50]. These abilities occur naturally in the last attention layer of DINOv2, where the attention maps computed between the CLS token and the spatial tokens align with relevant objects within the image (see Fig. 2). Despite the remarkable results observed on image-only tasks, DINOv2 lacks a solid bridge with natural language, making it impossible to directly compute the similarities with the text features, as expressed in Eq. 1.
While DINOv2 and CLIP embedding spaces are traditionally thought as being uncorrelated spaces, we show that the CLIP textual embedding space can be projected into the DINOv2 space through a learnable nonlinear warping.
@@ 3.2 Augmenting DINO with Semantics
Specifically, we first extract the $N$ attention maps $A_{i}\in\mathbb{R}^{\frac{H}{P}\times\frac{W}{P}}$ (one for each of the i=1,,Ni=1,...,N heads) which DINOv2 computes between the CLS vector and its patch features from the last layer. One of the key features of DINOv2 is that each $A_{i}$ highlights different semantic regions within the image. For each attention map $A_{i}$, we compute a visual embedding $v^{A_{i}}\in\mathbb{R}^{D_{v}}$ as a weighted average of the dense feature map $v$, emphasizing the spatial areas that $A_{i}$ highlights. We then calculate the cosine similarity between each $v^{A_{i}}$ and the projected text embedding $\psi(t)$, resulting in $N$ similarity scores. Formally, the cosine similarity score between a head and the text embedding is defined as
[…]
$\text{with }v^{A_{i}}$ $=\sum_{h,w}v_{[h,w]}\text{softmax}(A_{i})_{[h,w]}.$ (5)
To obtain the most relevant score for alignment, we apply a selection function over the similarity scores obtained for different heads. In particular, we choose the maximum similarity $\max_{i=1,...,N}\text{sim}(v^{A_{i}},t)$ score across all heads, therefore promoting a robust alignment between textual and visual representations that adapts to the most salient visual features corresponding to the text query.
@@
$\mathcal{L}_{\text{InfoNCE}}=-\frac{1}{2B}\sum_{i=1}^{B}\log\frac{\exp(\text{sim}(\tilde{v}_{i},t_{i}))}{\sum_{j=1}^{B}\exp(\text{sim}(\tilde{v}_{j},t_{i}))}$
$-\frac{1}{2B}\sum_{i=1}^{B}\log\frac{\exp(\text{sim}(\tilde{v}_{i},t_{i}))}{\sum_{j=1}^{B}\exp(\text{sim}(\tilde{v}_{i},t_{j}))}.$
@@ 3.3 Identifying Background Regions
and normalize $\mathcal{F}$ by applying a softmax normalization over both spatial axes, and linearly re-projecting its values in the range $\min_{j,h,w}\mathcal{S}(I,T_{j})_{[h,w]},\max_{j,h,w}\mathcal{S}(I,T_{j})_{[h,w]}]$, where $\mathcal{S}(\cdot)$ is the similarity map defined in Eq. 1. We exploit the resulting normalized average attention per category to shape the similarity map by activating the foreground region and deactivating the background. The resulting shaped similarity map $\overline{\mathcal{S}}\in\mathbb{R}^{H\times W}$ is defined as
@@
where $\lambda$ is a hyperparameter representing the relevance of the background shaping in computing the segmentation masks. The background mask is then identified as the collection of pixels for which the shaped similarity map is lower than a threshold across all semantic categories.
@@
Model Visual Encoder Frozen V20 C59 Stuff City ADE V21 C60 Object Avg V20 C59 Stuff City ADE V21 C60 Object Avg
@@
GroupViT [54] Custom ViT 79.7 23.4 15.3 11.1 9.2 50.4 18.7 27.5 29.4 - - - - - - - - -
ReCo [41] CLIP 57.7 22.3 14.8 21.1 11.2 25.1 19.9 15.7 23.5 - - - - - - - - -
TCL [9] CLIP 77.5 30.3 19.6 23.1 14.9 51.2 24.3 30.4 33.9 - - - - - - - - -
SILC [36] Custom ViT 77.5 31.6 20.8 26.9 19.3 - - - - - - - - - - - - -
MaskCLIP [63] CLIP 74.9 26.4 16.4 12.6 9.8 38.8 23.6 20.6 27.9 29.4 12.4 8.8 11.5 7.2 23.3 11.7 7.2 13.9
CLIP-DIY [51] CLIP+DINO 79.7 19.8 13.3 11.6 9.9 59.9 19.7 31.0 30.6 - - - - - - - - -
SCLIP [47] CLIP 80.4 34.2 22.4 32.2 16.1 59.1 30.4 30.5 38.2 70.6 25.2 17.6 21.3 10.9 44.0 22.3 26.9 29.9
CLIP-DINOiser [52] CLIP 80.9 35.9 24.6 31.1 20.0 62.1 32.4 34.8 40.2 - - - - - - - - -
ClearCLIP [25] CLIP 80.9 35.9 23.9 30.0 16.7 51.8 32.6 33.0 38.1 80.0 29.6 19.9 27.9 15.0 - - - -
NACLIP [18] CLIP 79.7 35.2 23.3 35.5 17.4 58.9 32.2 33.2 39.4 78.7 32.1 21.4 31.4 17.3 52.2 28.7 29.9 36.5
dino.txt [21] DINOv2(reg) - - - - - - - - - 62.1 30.9 20.9 32.1 20.6 - - - -
FreeDA [3] DINOv2 77.1 37.1 24.9 34.0 19.5 51.7 32.6 24.4 37.7 71.8 35.4 24.2 32.3 19.4 44.9 31.1 24.6 35.5
FreeDA [3] CLIP+DINOv2 84.3 39.7 25.7 34.1 20.8 51.8 35.3 36.3 41.0 85.7 39.7 26.3 33.6 21.4 44.1 34.8 33.9 39.9
ProxyCLIP [26] CLIP+DINOv2(reg) 83.0 37.2 25.4 33.9 19.7 58.6 33.8 37.4 41.1 85.2 36.2 24.6 35.2 21.6 56.6 33.0 36.7 41.1
ProxyCLIP [26] CLIP+DINO 80.3 39.1 26.5 38.1 20.2 61.3 35.3 37.5 42.3 83.2 37.7 25.6 40.1 22.6 60.6 34.5 39.2 42.9
Talk2DINO (Ours) DINOv2(reg) 87.1 39.8 28.1 36.6 21.1 61.5 35.1 41.0 43.8 87.1 39.1 27.0 35.8 21.1 60.1 34.2 37.6 42.8
@@
GroupViT [54] Custom ViT 81.5 23.8 15.4 11.6 9.4 51.1 19.0 27.9 30.0 - - - - - - - - -
ReCo [41] CLIP 62.4 24.7 16.3 22.8 12.4 27.2 21.9 17.3 25.6 - - - - - - - - -
TCL [9] CLIP 83.2 33.9 22.4 24.0 17.1 55.0 30.4 31.6 37.2 - - - - - - - - -
MaskCLIP [63] CLIP 72.1 25.3 15.1 11.2 9.0 37.2 22.6 18.9 26.4 - - - - - - - - -
SCLIP [47] CLIP 83.5 36.1 23.9 34.1 17.8 61.7 31.5 32.1 40.1 76.3 27.4 18.7 23.9 11.8 47.8 23.8 26.9 32.1
CLIP-DINOiser [52] CLIP 81.5 37.1 25.3 31.5 20.6 64.6 33.5 36.1 41.3 - - - - - - - - -
LaVG [22] CLIP+DINO 82.5 34.7 23.2 26.2 15.8 62.1 31.6 34.2 38.3 - - - - - - - - -
NACLIP [18] CLIP 83.0 38.4 25.7 38.3 19.1 64.1 35.0 36.2 42.5 84.5 36.4 24.6 37.1 19.6 57.9 36.4 34.6 41.4
FreeDA [3] DINOv2 79.5 40.2 27.1 34.4 20.9 52.0 35.2 25.8 39.4 75.2 39.0 27.0 33.1 21.3 45.3 34.3 26.7 37.7
FreeDA [3] CLIP+DINOv2 85.2 42.1 27.0 33.8 21.8 51.8 37.4 38.6 42.2 87.1 42.4 28.1 33.8 22.6 55.4 37.1 36.1 42.8
ProxyCLIP [26] CLIP+DINOv2(reg) 83.1 38.9 26.6 35.4 20.3 62.0 35.2 38.7 42.5 85.8 37.6 25.6 37.5 22.5 59.4 34.6 39.0 42.8
ProxyCLIP [26] CLIP+DINO 80.3 39.4 26.9 38.6 20.2 60.8 35.3 37.2 42.3 83.2 38.0 26.2 41.0 22.6 60.7 34.7 39.4 43.2
Talk2DINO (Ours) DINOv2(reg) 88.5 42.4 30.2 38.1 22.5 65.8 37.7 45.1 46.3 89.8 42.7 29.6 38.4 22.9 66.1 37.3 42.3 46.1
Table 1: Comparison with unsupervised OVS models on Pascal VOC [15], Pascal Context [34], COCO Stuff [7], COCO Object [7], Cityscapes [11], and ADE20K [61, 62]. For each method, we specify the visual backbone used, along with whether it is frozen or fine-tuned. We report both the variants with and without background for Pascal VOC (V21 and V20) and Pascal Context (C60 and C59). Best results with and without mask refinement are highlighted in bold, overall best results are underlined.